# RePrompt: Planning by Automatic Prompt Engineering for Large Language Models Agents

### v1

@@ 4.1 Experiment Settings
In the experiments, the stopping criteria and other hyperparameter settings for each domain will be the same as in the original environments. Without specifications, we use a temperature $=0$, and a seed of $42$. All results are tested on GPT-4-turbo-1106-preview, given its strong general capability (Chiang et al., 2024). To help reproducibility, we provide all the optimized prompts generated by RePrompt in the appendix.
To fully test the capability of our algorithm in different scenarios, we choose two environments, PDDL generation (Guan et al., 2023), and Travel Planner (Xie et al., 2024). The PDDL generation task provides accurate but expensive feedback and challenges the exact translation capability of the LLMs, which is necessary for LLMs to be further able to write correct code. The TravelPlanner environment, on the other hand, provides cheap but not accurate feedback through Reflexion without knowing ground-truth information. Travel Planner also provides tools to be used to query the cost information in the database and challenges the reasoning capability by asking for direct generating solutions. Furthermore, in TravelPlanner, we are testing RePrompt with Reflexion, which further includes the thought-action-observation steps rather than a standard step-by-step instruction in the PDDL where each step is an intermediate step that could be helpful for guiding the generation for final results. Given the different types of feedback, the purpose of our RePrompt also change: in PDDL, our RePrompt serves to improve the zero-shot generation performance that is used to reduce the cost of generating feedback, while in TravelPlanner, RePrompt is used to help guide the LLM to take all the important steps of concern in all scenarios and reduce the potential failure caused by incomplete thoughts.
@@ 4.2 PDDL Generation
We first test the Planning Definition and Domain Language (PDDL) generation task (Guan et al., 2023). Given a natural language description of a PDDL instance, the job is to define the precondition and the effects of the actions in PDDL. Specifically, we consider the very first step of constructing the model and do not further consider the later correction phase and PDDL translation phase.  [^1] At the time of submission of our paper, the evaluation phase is missing in the official Github repository, and we are not able to compare the success rate in those phases in a fair manner. After generating the preconditions and effects of the actions, human domain experts are introduced to check whether the generation is correct. In this paper, we not only have human domain experts but also use another LLM as a separate checker to verify whether any of the errors that appear in the results given by the prompt in the original paper (Guan et al., 2023), which are released together with the code of it, have also appeared in our result.
@@
In this experiment, we use the generated result and the annotation from human experts of the prompt from the previous paper Guan et al. (2023) in "Tyreworld" as the chat history used in the RePrompt training set, and update it for one epoch to get the updated prompt. Here, because the feedback is provided by domain experts, it is accurate but expensive, so multiple rounds of iterations are not feasible, and so we choose to only train RePrompt for one epoch and greatly reduce the need for extra annotations. As shown in Table. 1, the prompt we get from RePrompt not only outperforms in the set that we are training on, i.e., the Tyreworld domain, but also generalizes to other related domains and improves the success rate there. Interestingly, we found that after changing the prompt, the prompt does not introduce any new errors, i.e., the errors the new prompt made are a subset of the errors made by the prompt in the original paper. With this subset of errors, fewer domain experts will be needed to give annotations, and make the whole PDDL translation process much faster.
Among the remaining errors, some of them are caused by missing common knowledge that is addressed in the description. For example, in the action "Empty a Vacuum Cleaner", the action description includes one sentence saying, " The trash can should be opened if it’s openable." This is common knowledge, and it has no additional information. However, in the current context, this sentence indicates one precondition. However, the PDDL precondition generated by the LLM has been omitted, which leads to one error. Similar problems also happen a few times and contribute to a significant number of errors that could be classified together.
@@ 4.3 Travel Planner
Next, we test on the sole-planning setting in TravelPlanner benchmark (Xie et al., 2024).  [^2] We have slightly changed some code to recognize some specific format in the automatic checking procedure. We guarantee that all the changes are for the outputs that have previously thrown an error in the checker program, but not the ones that have already been automatically rejected. While this does not necessarily lead to a higher delivery rate, the delivery rate may show a small improvement compared to the original code, and it will not affect other metrics In this benchmark, the LLMs are required to provide a concrete day-to-day plan, including where they should stay, eat, and how they should travel, and satisfying both commonsense constraints like reasonable city routes and budget constraints. While there are some breakdowns of what specific kind of constraints the plan does not satisfy, the primary metric that is used for comparing different methods is the final pass rate. It needs to be addressed that in this benchmark, the evaluation is done after the act loop is done, separately with a ground-truth checker rather than directly to the feedback loop in Reflexion, and thus, the feedback in the chat history used by RePrompt does not actually involve any human interference or oracles on what is the correct answer and what is the list of constraints. This allows us to train on a subset of the test set without worrying about leaking any extra oracle information to the model. Because of this, we choose to report results on the validation set of 180 data points instead of results on the larger test set in order to save API costs for our model.  [^3] With the 180 validation data points tested by Reflexion strategy on sole-planning, the cost to finish one generation is \$300-800 depending on what exact reflection is provided to the model. The full test set includes 1000 data points, and the cost will grow around linearly in the number of data points tested. Because RePrompt is based on further collecting data, we choose a subset of 10 data points in the validation set as our training dataset.
@@
Table 2: Results on Travel Planner Benchmark. The best results are marked in bold. The delivery rate, commonsense pass rate, and hard constraint pass rate overall contribute to the final pass rate, which is the main metric in this table. Because we are "training" on part of the data, but not using any additional ground-truth information, we split the results into overall final pass rate and training set final pass rate. All the metrics in the table are better when the numbers are larger.
@@
As shown in Table. 2, the prompt generated by 5 epochs of training of RePrompt is better than the pure Reflexion result in the final pass rate. In the optimization procedure, unlike the PDDL environment, the optimized prompt after one epoch does not show any benefit. This is because, as we discussed earlier, the prompt after the first epoch will only include one additional suggestion, which is about looking into the budget constraint in this case. No matter what this round of updates is, it is something summarized from the thoughts provided by the ReAct scheme and something the iteration loop of generating a final plan has often already noticed and addressed. The optimized prompt after 5 epochs helps LLMs perform better in both the data set we used for training and the other data not included in the optimization process. This shows the generalizability of the prompt we get through the process.
Further looking into where we have gotten our improvement on the final pass rate, we found that our method helps in improving the macro commonsense pass rate (shown in Table 2), which is the major bottleneck in the current stage of using LLM for travel planning. Further breakdown, we found that the prompt has significantly improved the pass rate on the so-called "reasonable city route" (not shown in Table 2, but one of the constraints tested by TravelPlanner (Xie et al., 2024)). This common sense is to ensure "Changes in cities during the trip must be reasonable" (Xie et al., 2024). It is common sense that the first round result without interaction can mostly be satisfied but easily forgotten after the interactions with feedback providers. Thus, this is one of the common sense constraints that LLMs can detect themselves in the process, and it is sometimes, but not often, covered in the thoughts provided in the feedback loop. Our RePrompt has caught and included this constraint in the prompt, so most iterations will have covered it. We believe this is one example that our algorithm has addressed the challenge of "Agents struggle to align their actions with their reasoning." mentioned in the Travelplanner paper (Xie et al., 2024).
To our surprise, we found that our method does not help stop "Agents producing hallucinatory answers due to information confusion.", and our agents still output the wrong flight number that is used for the wrong leg of the flight, and output the same flight number for both departure and arrival flight. While this theoretically could be addressed by the feedback provider, it seems that such feedback is never provided and addressed in the loop, and our RePrompt training loop cannot catch such an error either. We are looking forward to further investigating how our RePrompt framework can serve as a further level of checker on simple hallucination errors. In which, one possibility is to combine scenario-specific information with chat history when we try to calculate the "loss".
5 Discussion
5.1 Error Analysis
In our experiments, our automatic prompt optimization process does not provide any guarantees of successfully generating a better prompt. In this section, we briefly enumerate a few errors that often occur and describe our methods to address them. We believe that these additional methods are neither necessary nor generally applicable to all domains, and we choose to leave them here just as an ad-hoc solution. We believe that all of these ad-hoc solutions can be automatically removed when the instruction-following capability of LLMs is further improved.
5.1.1 Incomplete Prompt
@@
Figure 2: An example output of the optimizer LLM that outputs a prompt template instead of a complete prompt. While it is technically correct and successfully added the additional instruction shown in blue, this output is not acceptable since it includes a template holder for examples, which is marked in red, and this output still needs post-processing with the original prompt to make the prompt complete. For simplicity, we have omitted part of the original prompts that is not changed, marked in green, and this part of the prompt can be found in the original paper (Guan et al., 2023).
There are many times that the prompt optimizer fails to generate the complete prompt. As shown in the complete prompt in the appendix, we have already added extensive instructions on generating a complete prompt rather than an incomplete one. As shown in Fig. 2, LLMs sometimes will output a prompt template where it requires a copy-paste done by the users to make the prompt complete. We found this happens often when the initial prompt is relatively long, and we believe the LLMs are trained to generate a shorter response if possible, but they failed to follow our instructions on generating a complete prompt. To solve this problem, we add an additional LLM to help fill in the template. We also provide the prompt of this LLM in the appendix. This additional LLM helps us to generate a complete prompt in the TravelPlanner domain. We do not use a rule-based fixer because the generated prompt template could be included in multiple symbols including but not limited to <> or {}. To relieve our workload, we choose to let LLM automatically recognize and replace them.
5.1.2 Incorrect Change by Accident
@@
In some domains, the output format can be similar to a more commonly used domain, and LLMs are misled to correct the prompt in certain parts. For example, in our PDDL domains, we ask the LLM to generate the preconditions of the actions rather than the actual PDDL file. In our experiment, we found that even though our prompt has explicitly required the LLM not to change any other part, especially the output format part, the updated prompt still sometimes changes the output format by mistake, specifically, changing the output of "Preconditions" in capital into "precondition" in smaller cases. To solve this problem, we leverage the feedback of the syntax checker. While the generated results could have some errors, the results should always be complete, and have the required syntax. And if our syntax parser that extracts the answer from the LLM output cannot find the word "Precondition", we know the prompt used is not correct, then we re-run RePrompt on the same step to generate a correct one. Because the fail rate with our current code is empirically less than $10\%$, this ad-hoc solution is enough.
5.2 Limitations
@@
6 Conclusion
In this paper, we have focused on optimizing the prompts used in LLM agents. We propose a new automatic prompt optimizer, RePrompt, which is based on the summarization of the interaction between LLM agents and feedback providers. Our experiments show that the LLM agents could benefit from an updated prompt in both 1-epoch settings and 5-epochs settings. We also discuss the current limitations of our paper and how future work can adjust them.
Appendix A Pseudo Code for RePrompt

### v2

@@ 4.1 Experiment Settings
To test the capability of our algorithm in different scenarios, we choose three environments, PDDL generation (Guan et al., 2023), TravelPlanner (Xie et al., 2024), and Meeting Planning from Natural Plan Zheng et al. (2024). The first two tasks are selected because their feedback generator is already included in the paper, and the reality is that both datasets are hard to conduct iterations with feedback generators. The Meeting Planning task is selected as another LLM agent task that differs from the existing ones.
The PDDL generation task provides accurate but expensive feedback and challenges the exact translation capability of the LLMs, which is necessary for LLMs to be further able to write correct code. The TravelPlanner environment, on the other hand, provides cheap but not accurate feedback through Reflexion without knowing ground-truth information. TravelPlanner also provides tools to be used to query the cost information in the database and challenges the reasoning capability by asking for direct generating solutions. Furthermore, in TravelPlanner, we are testing RePrompt with Reflexion, which further includes the thought-action-observation steps rather than a standard step-by-step instruction in the PDDL where each step is an intermediate step that could be helpful for guiding the generation for final results. For Natural Plan, we test the performance of deepseek-R1 et al. (2025), an open-source model that provides its complete long chain-of-thought, and we directly use its own thinking part in the output as the feedback as recent models like deepseek-R1 already included self-evaluation and self-correction in its outputs. Given the different types of feedback, the purpose of our RePrompt also change: in PDDL, our RePrompt serves to improve the generation performance without any iteration between the LLM actor and the feedback generator that is used to reduce the cost of generating feedback; in TravelPlanner, RePrompt is used to help guide the LLM to take all the important steps of concern in all scenarios and reduce the potential failures; in Meeting Planning, RePrompt is used to reduce the unnecessary trail happens in the thinking process of the LLM, and also reduce the randomness on the steps in the thoughts.
In the experiments, the stopping criteria and other hyperparameter settings for each domain will be the same as in the original environments. For PDDL generation and TravelPlanner, we use a temperature $=0$, a seed of $42$, and results are tested on GPT-4-turbo-1106-preview. To help reproducibility, we provide all the optimized prompts generated by RePrompt in Appendix E.
@@ 4.2 PDDL Generation
We first test the Planning Domain Definition Language (PDDL) generation task (Guan et al., 2023). Given a natural language description of a PDDL instance, the job is to define the precondition and the effects of the actions in PDDL. Specifically, we consider the very first step of constructing the model and do not further consider the later correction phase and PDDL translation phase.  [^1] At the time of submission of our paper, the evaluation phase is missing in the official Github repository, and we are not able to compare the success rate in those phases in a fair manner. After generating the preconditions and effects of the actions, human domain experts are introduced to check whether the generation is correct. In this paper, we not only have human domain experts but also use another LLM as a separate checker to verify whether any of the errors that appear in the results given by the prompt in the original paper (Guan et al., 2023), which are released together with the code of it, have also appeared in our result.
In this experiment, we use the generated result and the annotation from human experts of the prompt from the previous paper Guan et al. (2023) in “Tyreworld" as the chat history used in the RePrompt training set, and update it for one epoch to get the updated prompt. Here, because the feedback is provided by domain experts, it is accurate but expensive, so multiple rounds of iterations are not feasible, and so we choose to only train RePrompt for one epoch and greatly reduce the need for extra annotations. As shown in Table. 1, the prompt we get from RePrompt not only outperforms in the set that we are training on, i.e., the Tyreworld domain, but also generalizes to other related domains and improves the success rate there. Interestingly, we found that after changing the prompt, the prompt does not introduce any new errors, i.e., the errors the new prompt made are a subset of the errors made by the prompt in the original paper. With this subset of errors, fewer domain experts will be needed to give annotations, and make the whole PDDL translation process much faster.
Among the remaining errors, some stem from the omission of common knowledge explicitly stated in the description. For instance, in the action "Empty a Vacuum Cleaner," the description includes the sentence: "The trash can should be opened if it’s openable." While this is a commonsense statement with no new information, in the given context, it implicitly defines a necessary precondition. However, the LLM-generated PDDL precondition fails to capture this, leading to errors. Similar issues occur multiple times, collectively accounting for a significant portion of the observed errors, which can be categorized under this broader pattern.
@@ 4.3 TravelPlanner
Next, we test on the sole-planning setting in TravelPlanner benchmark (Xie et al., 2024). In this benchmark, the LLMs are required to provide a concrete day-to-day plan, including where they should stay, eat, and how they should travel, and satisfying both commonsense constraints like reasonable city routes and budget constraints. While there are some breakdowns of what specific kind of constraints the plan does not satisfy, the primary metric that is used for comparing different methods is the final pass rate. It needs to be addressed that in this benchmark, the evaluation is done after the act loop is done, separately with a ground-truth checker rather than directly to the feedback loop in Reflexion, and thus, the feedback in the chat history used by RePrompt does not actually involve any human interference or oracles on what is the correct answer and what is the list of constraints. This allows us to train on a subset of the test set without worrying about leaking any extra oracle information to the model. Because of this, we choose to report results on the validation set of 180 data points instead of results on the larger test set in order to save API costs. And because RePrompt is based on further collecting data, we choose a subset of 10 data points in the validation set as our training dataset. We report the same group of metrics defined in the original paper.
@@
Table 2: Results on TravelPlanner Benchmark. The best results are marked in bold. The delivery rate, commonsense pass rate, and hard constraint pass rate overall contribute to the final pass rate, which is the main metric in this table. Because we are “training" on part of the data, but not using any additional ground-truth information, we split the results into overall final pass rate and training set final pass rate. All the metrics in the table are better when the numbers are larger.
@@
As shown in Table. 2, the prompt generated by 5 epochs of training of RePrompt is better than the pure Reflexion Shinn et al. (2023) result in the final pass rate, and also outperform PromptAgent Wang et al. (2024) as an example of previous automatic prompt engineering work that was primarily designed for single-round question-answering task  [^2] Because their design was not suitable for Travelplanner, we have made necessary adaptations to make it testable in this case. We provide more details in Appendix B. . In the optimization procedure, unlike the PDDL environment, the optimized prompt after one epoch does not show any benefit. This is because, as we discussed earlier, the prompt after the first epoch will only include one additional suggestion, which is about looking into the budget constraint in this case. No matter what this round of updates is, it is something summarized from the thoughts provided by the ReAct scheme and something the iteration loop of generating a final plan has often already noticed and addressed. The optimized prompt after 5 epochs helps LLMs perform better in both the data set we used for training and the other data not included in the optimization process. This shows the generalizability of the prompt we get through the process. We observe that our baseline, PromptAgent, tends to change an extensive amount on the original prompt, and became even worse on the final test.
To better understand the source of our improvement in final pass rate, we analyzed its impact on macro commonsense pass rate—a key bottleneck in leveraging LLMs for TravelPlanner. As shown in Table 2, our method substantially enhances performance in this metric. Further analysis reveals that our prompt significantly improves the pass rate for the so-called "reasonable city route" constraint, a critical aspect tested by TravelPlanner (though not explicitly shown in Table 2). This common sense is to ensure “Changes in cities during the trip must be reasonable". This represents a commonsense constraint that LLMs can recognize during the process. However, while such constraints occasionally appear in the feedback loop’s reasoning, they are not consistently addressed. Our RePrompt framework successfully identifies and integrates this constraint into the prompt, ensuring that it is accounted for in most iterations. We believe this is one example that our algorithm has addressed the challenge of “Agents struggle to align their actions with their reasoning." mentioned in the original paper Xie et al. (2024). However, for our baseline PromptAgent, the extensive changes have made the LLMs fail to follow the desired format, leading to more useless steps and more failures.
Surprisingly, our method does not mitigate the issue of agents producing hallucinations due to information confusion. Specifically, our agents continue to generate incorrect flight numbers—assigning them to the wrong flight leg or erroneously using the same flight number for both departure and return flights. While, in theory, such errors could be corrected by the feedback generator, we observe that this feedback is rarely generated or incorporated into the prompt optimization loop. Moreover, our RePrompt training loop fails to detect these errors. Moving forward, we aim to explore how our RePrompt framework can serve as an additional verification layer for simple hallucination errors. One potential approach is incorporating scenario-specific information and chat history when computing the “loss."
4.4 Ablation Study
# Training Samples Final Pass Rate
1 1.22
2 2.44
5 3.44
10 3.89
25 2.00
Table 3: Number of Training Samples and Final Pass Rate for RePrompt on TravelPlanner.
Next, we provide an ablation study on the sampling size and the number of iterations. To fix the total budget spent on training the same, we fix a multiplication of the number of training samples times training epoch to 50. So, with more training samples, one will have fewer training epochs, which might lead to returning a prompt before its convergence. As a special case, when training samples are equal to 1, our method is a vanilla prompt optimizer without a summarizer that is done per scenario. We show our results in Table. 3. We found that selecting an appropriate number of training samples—balancing diversity with the number of epochs- can significantly enhance the performance of RePrompt. This aligns with our intuition, as RePrompt can be viewed as a form of gradient descent.
4.5 Meeting Planning
# Prompt Final Pass Rate (%)
Original 21.00
RePrompt 26.00
Table 4: Results on Meeting Planning from Natural Plan benchmark. Results are based on Deepseek-R1.
Lastly, we evaluate our algorithm on the Meeting Planning task, which involves scheduling meetings with friends while considering availability and travel-time constraints, aiming to maximize the total number of successful meetings. A key challenge of this task is that not every meeting can be scheduled for every problem instance, making it impossible to determine whether an optimal solution has been achieved. In real-world applications, this uncertainty complicates the feedback loop, as accurate assessments of solution quality are inherently difficult. Our approach leverages the think section of DeepSeek-R1, presenting a unique challenge for RePrompt: effectively extracting useful insights from long and unstructured chat histories. To ensure compatibility with the R1 model, we made specific modifications to the dataset, detailed in Appendix B. For training, we use the first nine data points, applying a batch size of three, and further evaluate performance on the first 100 data points.  [^3] We do not use the complete dataset mainly due to the limited availability of the Deepseek-R1 when the experiment is conducted.
The results, presented in Table 4, show that despite the presence of highly noisy thinking logs, RePrompt still improves the original prompt. This demonstrates RePrompt’s adaptability to the latest models, highlighting its ability to improve performance without relying on a specific type of feedback information.
5 Error Analysis
In our experiments, our automatic prompt optimization process does not guarantee the successful generation of an improved prompt. In this section, we outline common errors that arise and describe the ad-hoc strategies we employ to mitigate them. While these methods are neither universally necessary nor broadly applicable across domains, we include them here as pragmatic solutions to specific issues. We anticipate that as LLMs’ instruction-following capabilities continue to advance, these ad-hoc interventions will become obsolete and can be seamlessly eliminated.
5.1 Incomplete Prompt
@@
Here are two examples from the classical BlocksWorld domain to demonstrate the output format.
@@
Figure 2: An example output of the optimizer LLM that outputs a prompt template instead of a complete prompt. While it is technically correct and successfully added the additional instruction shown in blue, this output is not acceptable since it includes a template holder for examples marked in red, and this output still needs post-processing with the original prompt to complete the prompt. For simplicity, we have omitted the parts of the original prompt that are not changed, marked in green, and this part of the prompt can be found in the original paper (Guan et al., 2023).
The prompt optimizer occasionally fails to generate a complete prompt. As detailed in Appendix D, we have incorporated explicit instructions to ensure completeness; however, as shown in Fig. 2, LLMs sometimes produce prompt templates that require manual copy-pasting by users to finalize the prompt. This issue arises in both our algorithm, RePrompt, and our baseline, PromptAgent. We observed that this failure is more frequent when the initial prompt is relatively long, likely because LLMs are trained to generate concise responses whenever possible, despite our instructions to output a fully formed prompt. To address this, we introduce an additional LLM to automatically complete the prompt template. This approach enables us to generate fully structured prompts in the TravelPlanner domain. We opted against a rule-based fixer, as the generated templates use a variety of delimiters—including but not limited to <>absent<>< > and —making it impractical to manually define exhaustive replacement rules. Instead, we rely on the LLM’s ability to recognize and complete these templates autonomously, reducing manual intervention and improving efficiency.
5.2 Incorrect Change by Accident
@@
In some domains, the output format can be similar to a more commonly used domain, and LLMs are misled to correct the prompt in certain parts. For example, in our PDDL domains, we ask the LLM to generate the preconditions of the actions rather than the actual PDDL file. In our experiments, we found that even though our prompt has explicitly required the LLM not to change the output format part of the prompt text, the updated prompt still sometimes changes the output format by mistake, specifically, changing the output of “Preconditions" in capital into “precondition" in smaller cases. To solve this problem, we leverage the feedback of the syntax checker. While the generated results could have some errors, the results should always be complete, and have the required syntax. And if our syntax parser that extracts the answer from the LLM output cannot find the word “Precondition", we know the prompt used is not correct, then we re-run RePrompt on the same step to generate a correct one. Because the fail rate with our current code is empirically less than $10\%$, this ad-hoc solution is enough.
6 Conclusion
In this paper, we have focused on optimizing the prompts used in LLM agents. We propose a new automatic prompt optimizer, RePrompt, which is based on the summarization of the interaction between LLM agents and feedback generators and optimizes the step instructions in the prompt. Although the performance is still limited by the natural capability of the feedback generation loop, our experiments across three challenging reasoning tasks show that LLM agents could benefit from an updated prompt, regardless of the type of feedback provided and without the need for a ground-truth checker to be included in the procedure.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgments
The National Science Foundation (NSF) supported the research under grant number 2112533: "NSF Artificial Intelligence Research Institute for Advances in Optimization (AI4OPT)".
Appendix A Limitations
@@
Appendix B Implementation Details
B.1 PromptAgent Baseline
PromptAgent Wang et al. [2024] was initially proposed as an APE algorithm for question-answering tasks, which made its design mostly use the fact of ground-truth feedback generator which tells at least whether the final prediction is correct or not. In our TravelPlanner settings, the environment does not even provide this boolean checker during the APE process.
To adapt the PromptAgent framework to our settings, we created an LLM-based checker, which provided exactly the same input as the plan generator and the plan just generated. The checker is called after the reflection is finished and submitted a plan, i.e., each node in the MCTS tree of PromptAgent is now a complete reflection run. We provide the prompt of the checker in Fig. 4. During testing, we compared the performance of the LLM-based checker with a ground-truth checker, which is only available under specific dataset test conditions. The accuracy of the LLM-based checker is 89%. Although the 89% accuracy might seem reasonable, it’s important to note that the success rate of a plan is less than 5%. This means that even a checker that consistently marks plans as incorrect would achieve better performance overall. We observe the same type of hallucinations as the ones in the plan generation, which is one of the main reasons that there are many false negatives in this checker. Meanwhile, we have briefly tested 5 of the false negatives with the latest GPT-4o, and we found that, surprisingly, they are all correct. However, to make a fair comparison and for the consistent of models used in the experiment, we are still using GPT-4-turbo as the model for the checker.
Additionally, we have made some necessary changes to the gradient descent prompt template to remove the requirement of labels, also known as the ground-truth answers, which in our case is also not provided.
To make the comparison between Reprompt and PromptAgent fair, we use the Lite version of PromptAgent to limit the number of iterations. However, even with PromptAgent-Lite, it is still about twice as expensive compared to RePrompt, which shows another advantage of our algorithm.
B.2 Meeting Planning Dataset
{mdframed}
[linecolor=black, linewidth=1pt] You are a format transformer. You will be provided with a meeting plan, your job is to transform it into a more strict output format that start with word "You" and then to more detailed activity. Here are some examples:
You start at Russian Hill at 9:00AM.
You travel to Marina District in 7 minutes and arrive at 9:07AM.
You wait until 3:45PM.
You meet James for 75 minutes from 3:45PM to 5:00PM.
Try to change only the location, time, and name in your output for the above examples. Begin your final answer with: "The formatted output is: ".
==== Here is the solution you need to transform:
Solution
Figure 3: The prompt used to transform the output from Deepseek-R1 from 0-shot prompt to the desired format. The prompt should be fed into an LLM, which in the experiment is Deepseek-V3 to transform the output to meet the desired format of Natural Plan.
The dataset is provided in the format of directly providing prompts in both 0-shot settings and 5-shot settings. However, the checker from the original repository only support a specific output format, that is inferred only in the 5-shot version of the prompt. Nevertheless, the official guideline of Deepseek-R1 has instructed users to always use 0-shot settings et al. [2025] as few-shot prompts always significantly degrade the performance. Considering all of these, we use the 0-shot prompt as the initial prompt, collect the results, and use Deepseek-v3 to transform the collected solution to the desired output format. It need to be noticed that this additional process may introduce certain responses that should be correct to go wrong. We provide the prompt of this transformation in Fig. 3.
Appendix C Pseudo Code for RePrompt